Two complementary angles are in a ratio of 1 to 4. find the measure of each angle. (what are complementary angles? what is there some?)
Answer by checkley79(3341)

COMPLEMENTARY ANGLES ADD UP TO 90 DEGREES.
X+Y=90
Y=90-X
1/4=X/Y
1/4=X/(90-X) CROSS MULTIPLY.
4X=90-X
4X+X=90
5X=90
X=90/5
X=18 DEGREES FOR THE SMALLER ANGLE.
18+Y=90
Y=90-18
Y=72 DEGREES FOR THE LARGER ANGLE.
PROOF:
1/4=18/72
72=72